# BitMark for Infinity: Watermarking Bitwise Autoregressive Image Generative Models

### v1

@@ Appendix A Additional Algorithms
  • interpolate: Bilinear scaling to target resolution.
  • Count: Action performed for the statistical test, which counts the number of bits which are part of the green list.
  • $\phi_{i}$ extra convolution operations to address the information loss (following [32]).
We present the pseudo code of the novel Bit-Flipper attack in Algorithm 3. To apply the attack, we have to first encode the image in to bits (line 1). During deconstructing the old watermarked image, we construct a new attacked image. We recursively iterate through all scales $K$, retrieve the bit representations of the respective scale (line 4) and count how many bits are part of the green list and the red list (line 5). Then we define the probability of flipping a bit based on the green fraction and the flip factor $\phi$. Next, we change the bits which are part of the green list based on the random probability $p$ (line 8, 9). We follow the standard encoding and decoding process, to iterate through all resolutions of the image (line 10-13). Finally, we return the decoded version of the attacked image.
Algorithm 3  Bit-Flipper Attack
Inputs: image $im$, green list $G$, red list $R$, image encoder $\mathcal{E}$, quantizer $\mathcal{Q}$, image decoder $\mathcal{D}$.
Hyperparameters: Flip factor $\phi$, steps $K$ (number of resolutions), resolutions $(h_{i},w_{i})_{i=1}^{K}$, the number of bits for resolution $i$ is $t_{i}$, $n$ the length of the bit vector, $|s|$ the total number of bit sequences.
@@
$img=\text{init()}$()
for $i=1,\dots,K$ do
       $u_{i}=\mathcal{Q}(\text{interpolate}(e,h_{i},w_{i}))$ $C_{i}$ = Count($(b_{1},\dots,b_{t_{i}})$, $G$) $p=(\frac{C_{i}}{|s_{i}|}-0.5)*\phi$ for $j=n,\dots,t_{i}$ do
             if $(b_{j-n},\dots,b_{j})\in G\land\text{random}()<p$ then
                   $b_{j}=\lnot b_{j}$Flip bit with probability $p$ if sequence is in green list}
Return: attacked image $\mathcal{D}(img)$
Appendix B Green and Red List Selection
Selection of Red and Green Lists. The re-encoding loss presented in Section 3.1 also motivates the choice of a small length of the green list $n$, as larger consecutive bit sequences are more likely to be interrupted. Additionally, this auto-encoder has the property of re-encoding images towards an equal frequency of 0’s and 1’s, so it is not efficiently possible to bias each bit towards either 0 or 1, making the choice of $n=1$ unpractical (see Table 6). Thus we leverage the second smallest possible choice, $n=2$, resulting in 6 possible choices to separate $G$ and $R$ by keeping $|G|=|R|$.
As we have the constraint to not bias towards 0 or 1, as this type of pattern is more likely removed by the auto-encoder, and given the constraint shown that the same prefix must not be in the same list twice as discussed in Section 3.2, the most effective $G$ for BitMark are $G=\{01,10\}$ and $G=\{00,11\}$. We choose $G=\{01,10\}$ throughout all experiments.
Changing the Length of the Bit Sequence. We additionally ablate the size of the bit sequence in Table 6. For each sequence size $n$, if both lists should have the same size, we choose $\frac{2^{n}}{2}$ out of $2^{n}$ possibilities without replacement, so there are $\binom{2^{n}}{2^{n-1}}$ possibilities of constructing the $G$ and $R$. For the 2-bit sequences, flipping the green and red list has only limited impact on the final z-score, thus for the 3-bit sequences we only analyze all unique choices, i.e., we do not interchange green and red list, leaving us with 35 distinct possible green and red lists. For $n=3$, two $G$ have a similar $z$ compared to the best 2-bit sequences. These choices have the same properties as discussed above, namely (1) they equally often bias towards 0 and 1 and (2) none of the sequences share the same prefix, i.e., the biasing is always effective.
@@ Appendix B Green and Red List Selection
$\{000,001,100,110\}$$\{010,011,101,111\}$0.05 49.92% 49.98%
$\{000,001,101,111\}$$\{010,011,100,110\}$0.26 49.92% 49.98%
$\{000,001,100,111\}$$\{010,011,101,110\}$0.37 49.92% 49.98%
$\{000,001,110,111\}$$\{010,011,100,101\}$0.55 49.92% 49.98%
$\{000,010,101,111\}$$\{001,011,100,110\}$0.65 33.07% 49.48%
$\{000,010,110,111\}$$\{001,011,100,101\}$0.85 33.07% 49.48%
$\{000,001,011,101\}$$\{010,100,110,111\}$1.45 58.19% 50.24%
$\{000,010,011,101\}$$\{001,100,110,111\}$12.98 40.92% 49.52%
$\{000,001,010,111\}$$\{011,100,101,110\}$13.39 44.09% 49.68%
Figure 3: Longer sequences are less consistent after re-encoding. We analyze the consistency of sequences of length $n$ after re-encoding over the whole image, following the setup of Figure 1.
Consistency in Longer Sequences. We analyze the overlap of bit sequences with growing size in Figure 3 which shows that with increased length of the sequence, the consistency of the pattern decreases. This has a direct impact on the choice of $n$, as it shows that larger $n$ lead to less overlap, decreasing the impact of BitMark which is already noticeable for $n=3$ (see Table 6).
Appendix C Experimental Setup
C.1 Hyperparameters
@@ C.2 Environment
Appendix D Natural Images
Dataset $z$ Green Fraction
Imagenet1k 0.6311 (1.966) 0.5005 (0.0017)
MS-COCO2014 0.6773 (1.735) 0.5006 (0.0015)
MS-COCO2014 (Infinity-2B generated) 0.3620 (1.641) 0.5003 (0.0014)
To ensure the validity of our hypothesis we compute the average number of members of the green list and $z$ for natural images from different datasets, as well as images generated for the non-watermarked Infinity model. Table 7 shows that indeed, the green fraction for non-watermarked images is 50%.
Appendix E Additional Experiments
E.1 Detailed Analysis of BitMark
Next, we analyze the low-level behavior of BitMark during image generation. Table 9 depicts information about the generation of images with Infinity, averaged over 10,000 images. It shows that the entropy is low in earlier scales, meaning the bits are chosen with high certainty as these depend more the input prompt, but large on later scales, which means sampling either of the possible values for the bit can lead to different, but both high-quality content. This insight aligns with the re-encoding loss, which keeps more bits consistent on earlier scales.
Table 10, depicts the low-level information of applying BitMark with $\delta=2$. Applying BitMark changes up to 37% of bits per scale, yet the re-encoding loss only increases by 5% compared to the non watermarked image. Table 8 further shows the correlation between different analyzed factors. As expected, there is a high correlation between changed bits and entropy, as due to the soft-biasing we mostly change bits which have low entropy.
Entropy Changed Bits $z$ Green Fraction
Changed Bits 0.999 - - -
$z$ 0.706 0.689 - -
Green Fraction 0.918 0.917 0.642 -
Re-Encoding Loss 0.926 0.931 0.475 0.751
@@ E.1 Detailed Analysis of BitMark
Watermarking in Early vs Late Scales. We provide both qualitative and quantitative analysis for the scales in the Infinity architecture. We observe that the model produces high-level information in the image during the initial scales, forming a coarse shape of the main visual components, such as the object and the background (see Figure 4. In larger scales, the model predicts the details of the image (see Figure 5). As shown in Table 11, earlier and late scales contribute differently to the robustness of BitMark. Earlier scales are more robust against like CtrlRegen, as CtrlRegen does not change the content of the image itself.
Figure 4: Visualization of different images on different scales. Here, the image on scale $i$ refers to the cumulative image obtained by adding up $1$ to $i$ scales. We visualize five sets of images: 1) Original images without watermarking, 2) Watermarked (last 3), where the watermarks are generated on scale 11 to 13, 3) Differences (last 3), which is the difference between the original and watermarked images (last 3), 4) Watermarked (all), where the watermarks are generated on all scales, 5) Differences (all), which is the difference between the original and watermarked images (all). Here, we visualize scales 1-7.
Figure 5: Visualization of different images on different scales. Here, the image on scale $i$ refers to the cumulative image obtained by adding up $1$ to $i$ scales. We visualize five sets of images: 1) Original images without watermarking, 2) Watermarked (last 3), where the watermarks are generated on scale 11 to 13, 3) Differences (last 3), which is the difference between the original and watermarked images (last 3), 4) Watermarked (all), where the watermarks are generated on all scales, 5) Differences (all), which is the difference between the original and watermarked images (all). Here, we visualize scales 8-13.
@@
E.2 Image Quality
We analyze the impact of applying BitMark with different $\delta$ on the image quality. Figure 6 shows that applying BitMark with $\delta\leq 3$ keeps the semantic structure and quality of the image the same, while changing small features, e.g., the color of the train in the first row. For larger watermarking strengths, the image quality starts to deteriorate, where artifacts become visible in the final image.
E.3 Timing
We present the average elapsed time (in seconds) for the generation of an image with Infinity, when the watermark is applied to all scales (resolutions) and when using the standard generation process (without a watermark). For hardware information see Section C.2. Table 12 shows that applying BitMark results in negligible overhead to image generation and allows for a fast detection in under 0.5 seconds.
@@ E.3 Timing
Setup Seconds/Image
BitMark 2.5817 (0.0063)
No watermark 2.3196 (0.0043)
Detection 0.4491 (0.0046)
E.4 Radioactivity
@@
Table 13 reports the radioactivity of BitMark for the class-conditional autoregressive models, namely, VAR [32], which operates in the multi-scale manner similarly to Infinity, and RAR [37], which generates tokens only on a single scale in an autoregressive fashion. We find that BitMark exhibits detectable traces of radioactivity for the class-to-image autoregressive models models with TPR>>much-greater-than$>$ >1%@FPR=1%, with slightly lower traces for RAR (probably due to its single scale generation) than VAR. A potential cause of the overall relatively lower radioactivity for VAR and RAR is stemming from the class-conditional setting, instead of the text-to-image as in Infinity.
E.5 Rotation
Figure 7: BitMark is robust against rotation within ±$\pm 20$ degrees. The TPR@1%FPR for different $\delta$ and rotation degrees.
We ablate the robustness of BitMark against different degrees of rotation and different watermarking strengths $\delta$. We find that with $\delta=2$ BitMark is robust for rotation degrees ±$\pm 30°$. Additionally we find an increase in TPR@1%FPR for a rotation of 180°.
Appendix F Analysis for the Watermarks in the Sand Attack
As shown by the qualitative analysis in Figure 8, the watermarks in the sand attack causes a significant loss of details and produces many perceptible artifacts, limiting the applicability of this attack.
Appendix G Analysis for the Bit-Flipper Attack

### v2

@@ Appendix C Application of BitMark on Single-Scale models
  • Interpolate: Scaling to target resolution.
  • Count: Counts the number of bits which are part of the green list.
Algorithm 3  Watermark Embedding
Inputs: autoregressive model $f_{\theta}$ with parameters $\theta$, green list $G$, red list $R$, image decoder $\mathcal{D}$.
Hyperparameters: the number of tokens $r$, constant $\delta$ added to the logits of the greengreen list bits, $n$ - the length of the bit sequences in the lists ($G$ and $R$), $m$ the number of bits per token.
for $t=1,\dots,r$ do
  for $j=n,\dots,m$ do
     $c=(t-1)\cdot m+j\text{ //counter}$$p_{c}=\text{Bias}(G,l_{c},\delta)$$s_{c}=\text{Softmax}(p_{c})$$b_{c}=\text{Sample}(s_{c})$
$u=(b_{1},\dots,b_{r\cdot m})$$e=\text{Lookup}(u_{i})$$im=\mathcal{D}(e)$
Return: watermarked image imim
Algorithm 4  Watermark Detection
Inputs: raw image $im$, green list $G$, red list $R$, image encoder $\mathcal{E}$, quantizer $\mathcal{Q}$
Hyperparameters: the number of tokens $r$, $n$ - the length of the bit vector.
@@
$C=0$
$u_{i}=\mathcal{Q}(e)$
$u_{i}=(b_{1},\dots,b_{r\cdot m})$
[…]
Appendix D Green and Red List Selection
Consistency in Longer Sequences. We analyze the overlap of bit sequences with growing size in Figure˜3 which shows that with increased length of the sequence, the consistency of the pattern decreases. This has a direct impact on the choice of $n$, as it shows that larger $n$ lead to less overlap, reducing the impact of BitMark which is already noticeable for $n=3$ (see Table˜7).
Selection of Red and Green Lists. The re-encoding loss presented in Section˜3.1 also motivates the choice of a small length of the green list $n$, as larger consecutive bit sequences are more likely to be interrupted. Additionally, this auto-encoder has the property of re-encoding images towards an equal frequency of 0’s and 1’s, so it is not efficiently possible to bias each bit towards either 0 or 1, making the choice of $n=1$ unpractical (see Table˜7). Thus we leverage the second smallest possible choice, $n=2$, resulting in 6 possible choices to separate $G$ and $R$ by keeping $|G|=|R|$.
As we have the constraint to not bias towards 0 or 1, as this type of pattern is more likely removed by the auto-encoder, and given the constraint shown that the same prefix must not be in the same list twice as discussed in Section˜3.2, the most effective $G$ for BitMark are $G=\{01,10\}$ and $G=\{00,11\}$. We choose $G=\{01,10\}$ throughout all experiments.
Figure 3: Longer sequences are less consistent after re-encoding. We analyze the consistency of sequences of length $n$ after re-encoding over the whole image, following the setup of Figure˜1.
Changing the Length of the Bit Sequence. We additionally ablate the size of the bit sequence in Table˜7. For each sequence size $n$, if both lists should have the same size, we choose $\frac{2^{n}}{2}$ out of $2^{n}$ possibilities without replacement, so there are $\binom{2^{n}}{2^{n-1}}$ possibilities of constructing the $G$ and $R$. For the 2-bit sequences, flipping the green and red list has only limited impact on the final z-score, thus for the 3-bit sequences we only analyze all unique choices, i.e., we do not interchange green and red list, leaving us with 35 distinct possible green and red lists. For $n=3$, two $G$ have a similar $z$ compared to the best 2-bit sequences. These choices have the same properties as discussed above, namely (1) they equally often bias towards 0 and 1 and (2) none of the sequences share the same prefix, i.e., the biasing is always effective.
@@ Appendix D Green and Red List Selection
$\{000,001,100,110\}$ $\{010,011,101,111\}$ -0.05 49.92% 49.98%
$\{000,001,101,111\}$ $\{010,011,100,110\}$ -0.26 49.92% 49.98%
$\{000,001,100,111\}$ $\{010,011,101,110\}$ -0.37 49.92% 49.98%
$\{000,001,110,111\}$ $\{010,011,100,101\}$ -0.55 49.92% 49.98%
$\{000,010,101,111\}$ $\{001,011,100,110\}$ -0.65 33.07% 49.48%
$\{000,010,110,111\}$ $\{001,011,100,101\}$ -0.85 33.07% 49.48%
$\{000,001,011,101\}$ $\{010,100,110,111\}$ -1.45 58.19% 50.24%
$\{000,010,011,101\}$ $\{001,100,110,111\}$ -12.98 40.92% 49.52%
$\{000,001,010,111\}$ $\{011,100,101,110\}$ -13.39 44.09% 49.68%
Appendix E Experimental Setup
E.1 Hyperparameters
@@ E.2 Environment
Appendix F Additional Experiments
In this Section we ablate the behavior of BitMark on the Infinity-2B model.
F.1 Scalewise Analysis of BitMark
Next, we analyze the scalewise impact of BitMark during image generation. Table˜8 depicts information about the generation of images with Infinity-2B, averaged over 10,000 images. It shows that the entropy is low in earlier scales, meaning the bits are chosen with high certainty as these depend more on the input prompt, but large on later scales, which means sampling either of the possible values for the bit can lead to different, but both high-quality content.
@@ F.1 Scalewise Analysis of BitMark
Table˜9, depicts using BitMark with $\delta=2$. Applying BitMark changes up to 37% of bits per scale, yet the re-encoding loss only increases by 5% compared to the non watermarked images. Table˜10 further shows the correlation between different analyzed factors. As expected, there is a high correlation between changed bits and entropy, as due to the soft-biasing we mostly change bits of low entropy.
Entropy Changed Bits $z$ Green Fraction
Changed Bits 0.999 - - -
$z$ 0.706 0.689 - -
Green Fraction 0.918 0.917 0.642 -
Re-Encoding Loss 0.926 0.931 0.475 0.751
We further observe that the model produces high-level information in the image during the initial scales, forming a coarse shape of the main visual components, such as the object and the background (see Figure˜5. In larger scales, the model predicts the details of the image (see Figure˜6). As shown in Table˜11, earlier and late scales contribute differently to the robustness of BitMark. Earlier scales are more robust against like CtrlRegen+, as CtrlRegen+ does not change the content of the image itself, but mostly high level features.
@@
F.2 Robustness
Table˜12 shows an ablation of the robustness of BitMark given different values of $\delta$. BitMark achieves a high TPR@1%FPR against most attacks even for a low watermarking strength $\delta=1$. While stronger robustness can be achieved with higher $\delta$ values, there is a direct tradeoff with the image quality, which we show qualitatively in Figure˜7.
Table 12: BitMark is robust against watermark removal attacks. We report the TPR@1%FPR (%) for the different attacks and baselines.
$\delta$ Conventional Attacks Reconstruction Attacks
None Noise Blur Color Crop Rotation JPEG SD2.1-VAE CtrlRegen+
1 100.0 97.8 99.2 99.3 68.5 14.5 100.0 100.0 61.9
2 100.0 99.6 99.9 99.8 98.8 20.1 100.0 100.0 91.6
3 100.0 99.7 99.9 99.9 99.8 23.8 100.0 100.0 97.0
4 100.0 99.9 99.9 100.0 99.9 29.2 100.0 100.0 99.0
5 100.0 100.0 100.0 99.9 100.0 34.3 100.0 100.0 99.8
Rotation.
Figure 4: BitMark is robust against rotation within $\pm 30$ degrees. The TPR@1%FPR for different $\delta$ and rotation degrees.
We ablate the robustness of BitMark against different degrees of rotation and different watermarking strengths $\delta$. We find that with $\delta=2$ BitMark is robust for rotation degrees $\pm 30$°. Additionally we find an increase in TPR@1%FPR for a rotation of 180°.
We further show the improved detection performance leveraging a rotation estimation tool [21] in Table˜13. After re-rotating the image, the detection performance increases from TPR@1%FPR=0.184 to TPR@1%FPR=0.983. In addition, we consider an even harder case: after the random rotation, we apply a central crop to the image with 50% cropping ratio, such that all the black corners during rotation are completely removed. We show the results in Table˜14.
Table 13: Impact of rotation correction on detection performance. Images are randomly rotated between 0° and 360°.
Setting AUC (%) Accuracy (%) TPR@1%FPR (%) Green Fraction (%)
No rotation correction 60.4 60.8 18.4 50.2
After rotation correction 99.5 99.2 98.3 52.0
Table 14: Impact of rotation correction and cropping on detection performance. Images are randomly rotated between 0° and 360° with 50% center crop.
Setting AUC (%) Accuracy (%) TPR@1%FPR (%) Green Fraction (%)
No rotation correction 59.8 58.9 9.9 50.1
After rotation correction 89.1 86.5 66.3 50.5
F.3 Image Quality
Table˜15 shows that our chosen biases $\delta$ does not negatively impact the image quality for Infinity-2B ($\delta=2$), Infinity-8B ($\delta=1.5$) and Instella IAR ($\delta=1.5$)
Additionally, we qualitatively analyze the impact of applying BitMark with different $\delta$ values. Figure˜7 shows that applying BitMark with $\delta\leq 3$ keeps the semantic structure and quality of the image the same, while changing small features, e.g., the color of the train in the first row. For larger watermarking strengths, the image quality starts to deteriorate, where artifacts become visible in the final image.
Table 15: BitMark does not decrease the image quality. Changes in FID, KID and CLIP relative to non-watermarked images.
Model $\Delta\mathrm{FID}$ $\Delta\mathrm{KID}$ $\Delta\mathrm{CLIP}$
Infinity 2B, $\delta=2$ -2.1169 -0.0016 0.1086
Infinity 8B, $\delta=1.5$ 0.5299 -0.0018 0.2792
Instella IAR, $\delta=1.5$ 1.1471 0.0004 -0.8361
F.4 Timing
@@
BitMark No watermark Detection
Seconds/Image 2.5817 (0.0063) 2.3196 (0.0043) 0.4491 (0.0046)
We present the average elapsed time (in seconds) for the generation of an image with Infinity, when the watermark is applied to all scales (resolutions) and when using the standard generation process (without a watermark). For hardware information refer to Section˜E.2. Table˜16 shows that BitMark results in negligible overhead to image generation and allows for a fast detection in under 0.5 seconds.
F.5 Radioactivity to IARs
@@
Table˜17 reports the radioactivity of BitMark for the class-conditional autoregressive models, namely, VAR [35], which operates in the multi-scale manner similarly to Infinity, and RAR [41], which generates tokens only on a single scale in an autoregressive fashion. We find that BitMark exhibits detectable traces of radioactivity for the class-to-image autoregressive models models with TPR>>>>1%@FPR=1%, with slightly lower traces for RAR (probably due to its single scale generation) than VAR. A potential cause of the overall relatively lower radioactivity for VAR and RAR is stemming from the class-conditional setting, instead of the text-to-image setting as in Infinity. +
F.6 Natural Images
To ensure the validity of our hypothesis we compute the average number of members of the green list and $z$ for natural images from different datasets, as well as images generated for the non-watermarked Infinity model. Table˜18 shows that indeed, the green fraction for non-watermarked images is 50%.
Dataset $z$ Green Fraction
Imagenet1k 0.6311 (1.966) 0.5005 (0.0017)
MS-COCO2014 0.6773 (1.735) 0.5006 (0.0015)
Infinity-2B generated 0.3620 (1.641) 0.5003 (0.0014)
We additionally show the FPR for different natural datasets (ImageNet-1k, LAION POP) and Infinity-2B generated images, given the 1%FPR threshold that we used for MS-COCO in Table˜19.
Table 19: False Positive Rate of choosing the MS-COCO 1%FPR threshold on different datasets. FPR measured on N=1000 non-watermarked samples.
Dataset FPR@1%MS-COCO
MS-COCO 0.01
ImageNet-1k 0.011
LAION POP 0.021
Infinity-2B Generated 0.0
Appendix G Generalization to Diffusion
Additionally to the bitwise autoregressive image modeling that we explore in Section˜4 we extend our BitMark to bitwise diffusion models, namely BiGR [12]. We employ two variations of BiGR, the L-256x256 with 24 bits per token and L-512x512 which leverages 32 bits per token. We set the Gumbel temperature to 0.01, the cfg to 2.5, the number of iterations as well as the diffusion timesteps to 10 each. We further notice that the detection performance for BiGR increases greatly if, before encoding, a slight amount of noise is added to the image.
In Table˜20 we show that BitMark has negligible impact on the quality of the generated images for both BiGR model sizes. We use 5,000 MS-COCO prompts to generate both BiGR clean and watermarked data. On these images we then compute image quality metrics FID, KID and CLIP. The difference in FID and KID score shows that the watermarked images have lower FID and KID, a sign of more similarity to the MS-COCO baseline images. The CLIP score only increases slightly, showing only minor impact of BitMark to the prompt following capabilities of the model.
Additionally we show the robustness of BitMark in Table˜21, it achieves significant TPR@1%FPR for all attacks. Especially against the advanced reconstruction attack BitMark achieves 75.2 and 69.2%TPR@1%FPR at resolution of $256\times 256$ and $512\times 512$, respectively. We note that, BiGR generates smaller images with only 256 or 1024 tokens, leading to a lower robustness compared to the Infinity-2B and 8B model.
Table 20: BitMark does not decrease the image quality. Changes in FID, KID and CLIP relative to non-watermarked images.
Model $\Delta\mathrm{FID}$ $\Delta\mathrm{KID}$ $\Delta\mathrm{CLIP}$
BiGR 256x256, $\delta=3$ -3.0304 -0.0020 0.0017
BiGR 512x512, $\delta=3$ -3.6660 -0.0024 0.1329
Table 21: Watermark robustness against removal attacks. We report the robustness (%TPR@1%FPR) for the different attacks. For BiGR, we added a slight amount of noise to all images before encoding.
Watermark Conventional Attacks Reconstruction Attacks
None Noise Blur Color Rotate Crop JPEG Vertical Horizontal VAE CtrlRegen+
BiGR $256\times 256$ ($\delta=3$) 97.0 54.3 94.2 16.1 55.3 49.7 95.1 61.3 47.8 95.5 75.2
BiGR $512\times 512$ ($\delta=3$) 99.3 57.4 96.4 34.2 65.1 69.4 97.6 75.4 77.2 98.2 69.2
Appendix H Analysis of the Advanced Attacks
H.1 Analysis for the Bit-Flipper Attack
We present the pseudo code of the novel Bit-Flipper attack for multi-scale models in Algorithm˜5. To apply the attack, we first encode the image into bits (line 1). During deconstructing the watermarked image, we construct a new attacked image. We recursively iterate through all scales $K$, retrieve the bit representations of the respective scale (line 4) and count how many bits are part of the green list and the red list (line 5). Then we define the probability of flipping a bit based on the green fraction and the flip factor $\phi$. Next, we change the bits which are part of the green list based on the random probability $p$ (line 8, 9). We follow the standard encoding and decoding process, to iterate through all resolutions of the image (line 10-13). Finally, we return the decoded version of the attacked image.
Algorithm 5  Bit-Flipper Attack
Inputs: Image $im$, green list $G$, red list $R$, image decoder $\mathcal{D}$.
Hyperparameters: Flip factor $\phi$, steps $K$ (number of resolutions), resolutions $(h_{i},w_{i})_{i=1}^{K}$, the number of bits for resolution $i$ is $t_{i}$, $n$ - the length of the bit vector, $|s|$ the total number of bit sequences.
$e=\mathcal{E}(im)$
$img=\text{init()}$
for $i=1,\dots,K$ do
[…]
     if $(b_{j-n},\dots,b_{j})\in G\land\text{random}()<p$ then
        $b_{j}=\lnot b_{j}$ {Flip bit with probability $p$ if sequence is in green list}
Return: attacked image $\mathcal{D}(img)$